# Heterogeneous Intertemporal Treatment Effects via Dynamic Panel Data Models

### v1

@@ Abstract
We study the identification and estimation of heterogeneous, intertemporal treatment effects (TE) when potential outcomes depend on past treatments. First, applying a dynamic panel data model to observed outcomes, we show that instrument-based GMM estimators, such as Arellano and Bond (1991), converge to a non-convex (negatively weighted) aggregate of TE plus non-vanishing trends. We then provide restrictions on sequential exchangeability (SE) of treatment and TE heterogeneity that reduce the GMM estimand to a convex (positively weighted) aggregate of TE. Second, we introduce an adjusted inverse-propensity-weighted (IPW) estimator for a new notion of average treatment effect (ATE) over past observed treatments. Third, we show that when potential outcomes are generated by dynamic panel data models with homogeneous TE, such GMM estimators converge to causal parameters (even when SE is generically violated without conditioning on individual fixed effects). Finally, we motivate SE and compare it with parallel trends (PT) in various settings with observational data (when treatments are dynamic, rational choices under learning) or experimental data (when treatments are sequentially randomized).
@@ 1 Introduction
Consider a panel data setting in which a researcher observes individual outcomes $Y_{it}$ and treatments $D_{it}$ over multiple periods $t=1,...,T$. At any given time $t$, potential outcomes depend on a full sequence of potential treatments in the past, i.e., $Y_{it}(d^{t})$ where $d^{t}\equiv(d_{s})_{s=0}^{t}$ with $d_{s}\in\{0,1\}$. The objective of causal inference in this context is to learn about certain moments of $Y_{it}(d^{t})$ using the joint distribution of $\{D_{it},Y_{it}\}_{t\leq T}$, where the observed outcomes are determined as $Y_{it}=Y_{it}(D_{i}^{t})$.
This paper considers the question of causal inference through the lens of (linear) panel data models that also include a lagged outcome as a regressor, e.g.,
@@
There are at least two reasons one might argue for the necessity of such a dynamic panel data model (DPDM). First, treatment may only be randomly assigned conditional on the past history. For example, when estimating the effect of a training program on earnings or employment outcomes, enrollment into the training program may depend on previous earnings and employment (Ashenfelter, 1978). In such cases, the lagged outcome is a necessary control for the exogeneity of treatment. Even though (1) appears to deviate from a static panel data model only by an additional control, the fact that this additional control is a lagged dependent variable raises concerns about the strict exogeneity assumption that is necessary for two-way fixed-effect (TWFE) regressions. Second, outcomes may exhibit state-dependence. For example, past employment may have a causal effect on present employment (Heckman, 1981). In such cases, the lagged outcome also has a causal effect and may furthermore mediate the effects of past treatment. A DPDM such as (1) allows researchers to disentangle the partial effects of $D_{it}$ from the state dependence in $Y_{it-1}$.
One of our main questions is: when and how can we use existing dynamic panel data estimators (DPDE) of observed outcomes, particularly the GMM/2SLS estimators of Anderson and Hsiao (1982) and Arellano and Bond (1991), to shed light on causal effects? These estimators use lagged regressors as instruments to consistently estimate $\beta,\gamma$ in the first-differenced transformation of (1):
@@
where $\Delta$ denotes the first-difference operator. To answer this question, we begin by deriving a general causal decomposition of the aforementioned DPDEs (Lemma 1). Implicit in this decomposition, and more generally in our goal of causal inference, is the possibility that (1) and (2) are not structural models of the outcome-generating process, for example because of the possibility of treatment effect heterogeneity. This attempt to link such instrumented first-difference estimators of a reduced-form model to interpretations in terms of aggregated heterogeneous treatment effects (TE) is one contribution of this paper.
@@
The preceding exercise bears resemblance to, but differs qualitatively from, the work of De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021) in decomposing and interpreting TWFE regressions. In a non-dynamic setting where potential outcomes are indexed only by contemporary treatment $Y_{it}(d_{t})$, they showed the TWFE estimand in static panel data models (a special case of (1) with $\gamma=0$) is a weighted average of heterogeneous TE $Y_{it}(1)-Y_{it}(0)$. [^1] The use of TWFE in static panel data models requires that $D_{it}$ be strict exogenous in order to identify $\beta$ and $\gamma$. However, the difference-in-differences approach in De Chaisemartin and d’Haultfoeuille (2020) does not impose that the true data-generating process satisfies the parametric, reduced-form $Y_{it}=\beta D_{it}+\theta_{t}+\alpha_{i}+\varepsilon_{it}$. Their focus is instead on relating the TWFE estimand to aggregated heterogeneous treatment effects under parallel trends; they also consider first-difference regressions under the same assumptions. This literature has also considered the role and the complications of covariates (Heckman, Ichimura, and Todd, 1997; Abadie, 2005; Caetano and Callaway, 2023), as well as proposing alternative estimators under parallel trends (PT) and other assumptions on the outcome or treatment processes, like staggered designs (e.g.,De Chaisemartin and d’Haultfoeuille (2020); Callaway and Sant’Anna (2021); Sun and Abraham (2021)). In comparison, we face different challenges, because the treatment effects in our setting vary with the full sequence of past potential treatments, and the estimator in question uses instruments designed to address the failure of strict exogeneity of $D_{it}$ and $Y_{it-1}$ in the DPDM (1). In the context of papers imposing variants of parallel trends, the setting of de Chaisemartin and D’Haultfoeuille (2024) is most similar in that it also allows for dynamic potential outcomes and non-staggered designs; in contrast, however, our approach is based on a set of assumptions different than parallel trends.
Building on our decomposition in Lemma 1, we derive sufficient conditions under which a 2SLS first-differenced estimator as in Anderson and Hsiao (1982) is guaranteed to be a positively weighted average of heterogeneous TE (Proposition 2). These conditions include the sequential exchangeability assumption of Robins (1986) — which posits that the present potential outcomes are (mean) independent from contemporary treatment, once conditioned on a full history of observed outcomes and treatments — as well as restrictions on the heterogeneity of TE. We extend and strengthen the result by introducing an adjusted inverse-(treament) propensity-weighted (IPW) estimator and show that a sample analogue is consistent for a particular average treatment effect. This new estimator dispenses with the restrictions on heterogeneous TE in Proposition 2; it only requires sequential exchangeability and some mild conditions on treatment assignment. (See Proposition 3.)
@@
By relying on an assumption of sequential exchangeability, our results relate closely to a broad suite of treatment effect estimators collectively referred to as $g$-methods that have been developed for causal inference in this setting [e.g., see Robins (1986, 1997) and Robins and Hernán (2008) for a survey]. Most closely related, Robins (1997) proposes an IPW estimator that identifies $E[Y_{it}(d^{t})]$ using the observed outcomes $Y_{it}$ in stratum $\{D_{i}^{t}=d^{t}\}$; these potential outcomes can in turn be used to identify history-dependent treatment effects such as $E[Y_{it}(d^{t-1},1)-Y_{it}(d^{t-1},0)]$. In contrast, our adjusted IPW estimator identifies an aggregate treatment effect $E[Y_{it}(D_{i}^{t-1},1)-Y_{it}(D_{i}^{t-1},0)]$ using the first-differenced outcomes $\Delta Y_{it}$ among the entire sample. We conjecture that such an aggregation of treatment effects may be useful both from the standpoint of estimation and interpretation. In particular, our target parameter provides a potentially useful summary causal effect in multi-period settings where the number of potential contrasts grows large.
@@
The sequential exchangeability condition has played a central role in our results above (i.e., causal interpretation of dynamic panel data regression, and the adjusted IPW). This naturally raise two additional questions: (a) Can causal inference be conducted using DPDEs without relying on sequential exchangeability? (b) Is sequential exchangeability feasible in observational data where treatments are endogenously determined by forward-looking individuals? We provide conditions under which the answers to both questions are indeed positive.
@@
To answer (a), Propositions 4 and 5 delineate a case where our previous DPDEs identify well-defined causal estimands, and yet sequential exchangeability is generically violated without conditioning on further unobservables. Specifically, we focus on a parametric case where the data-generating process of the latent potential outcomes $Y_{it}(d^{t})$ follows a linear DPDM:
$Y_{it}(d^{t})=\beta^{\ast}d_{t}+\gamma^{\ast}Y_{it-1}(d^{t-1})+\theta_{t}^{\ast}+\alpha_{i}^{\ast}+\varepsilon_{it}^{\ast}(d_{t})\text{.}$ (3)
@@
In this case, the observed outcomes $Y_{it}=Y_{it}(D_{i}^{t})$ are consistent with the reduced form DPDM in (1), where the coefficients coincide with constant causal effects (i.e., $\beta=\beta^{\ast},\gamma=\gamma^{\ast}$ as noted in Observation 2). A further assumption of conditional sequential exogeneity condition on the structural errors $\varepsilon_{it}^{\ast}(\cdot)$ guarantee consistency of our standard DPDEs but not sequential exchangeability, because the exogeneity between contemporaneous outcomes and treatments now only holds after conditioning on unobserved fixed effects. Hence, this allows for endogenous selection on such unobservables.
@@
To investigate (b), we provide, in the context of observational data, conditions under which sequential exchangeability holds. This provides critical guidance for empiricists applying g-methods to observational data, especially when treatment choices are most likely dependent upon past outcomes. For example, when estimating the effect of a training program on earnings or employment outcomes, enrollment into the training program may depend on previous earnings and employment Ashenfelter (1978).
@@
To this end, we consider sequential exchangeability in a model of dynamic choices that serves as an anchor for structurally motivating the conditional and unconditional versions of sequential exchangeability. A particular focus of the model is on the possibility of learning, which is compatible with the nature of sequential exchangeability assumptions that condition (like a learning decision-maker) on the history of observed outcomes. The structural model provides a coherent dynamic decision problem of an economic agent where preferences, beliefs and decision rules are specified and shows the kinds of restrictions that are required, especially on the distribution of prior beliefs, to yield the exogeneity assumptions we use. Thus, our work connects to recent work on identification of learning models in nonparametric settings (e.g., (Bunting, Diegert, and Maurel, 2022)). Also, we connect a conditional independence assumption in the learning model to common parallel trends assumptions and show how sequential random assignment of treatment can be compatible with sequential exogeneity as expected but not parallel trends.
@@
Our work also relates to a broader agenda relating sequential exchangeability and parallel trends (e.g., Renson, Hudgens, Keil, Zivich, and Aiello (2023)) and the corresponding estimators, for example through “bracketing” relationships on the target parameter (Angrist and Pischke, 2008; Ding and Li, 2019). Additionally, Han (2021) studies the identification of dynamic treatment effects beyond sequential exchangeability, namely under a dynamic version of rank similarity. In complementary work, Kim (2023) studies identification of a dynamic panel data model when treatment is staggered and potential outcomes are non-negative, statically indexed, and take a multiplicative form; in this case, he obtains identification results for ratios of treatment effects under sequential exchangeability conditional on persistent heterogeneity.
The rest of the paper is organized as follows. Section 2 introduces the dynamic potential outcome framework. Section 3 studies the causal interpretation of DPDEs in the econometrics literature. Section 4 studies sequential exchangeability in the context of a model of dynamic choice with learning and illustrates the compatibility of parallel trends assumptions with conditional independence restrictions within the learning model and random assignment of treatments. Section 5 concludes.
@@ 2 A Dynamic Framework for Potential Outcomes
The researcher observes treatment $D_{it}$ with realized outcomes $Y_{it}$ for individual units $i$ in time periods $t=0,1,\dots,T$. There is a pre-treatment period $t=0$ with an initial realized outcome $Y_{i0}$ and $\Pr\{D_{i0}=0\}=1$. We take $Y_{i0}$ as given and unmodeled, thus avoiding the initial conditions problem (Heckman, 1981). Throughout what follows, we consider models in which potential outcomes in period $t$ are functions only of an individual’s own treatment history through period $t$.
@@
Potential outcomes for individual unit $i$ in period $t$ are indexed only by $i$’s own treatment through that time period, i.e., $Y_{it}(d_{i}^{t})$.
Assumption 1 rules out spillovers across individual units as well as anticipation of future treatments affecting present potential outcomes. Realized outcomes equal potential outcomes evaluated at realized treatments, $Y_{it}=Y_{it}(D_{i}^{t})$. We adopt notation $Y_{i}^{t}\equiv(Y_{i0},Y_{i1},\dots,Y_{it})$ for realized random (outcome) vectors, $Y_{i}^{t}(d^{t})=(Y_{i0},Y_{i1}(d^{1}),\dots,Y_{it}(d^{t}))$ for potential random vectors, and $Y_{i}^{t}(\cdot)$ for the collection of potential random vectors through period $t$. Throughout, we restrict to realized and potential treatment vectors $D_{i}^{t}$ and $d_{i}^{t}$ with degenerate initial realization $D_{i0}=d_{i0}=0$, which we suppress where convenient.
The main concern that motivates the inclusion of lagged outcomes as controls in a dynamic panel data model (DPDM) for observed outcomes in (1) is that the treatment is (mean) independent of potential outcomes only after conditioning on the observed history of past treatment and outcome. The assumption below formalizes such conditional independence. For subsequent references, we include both a full and mean version of the assumption. We drop individual subscripts $i$ to simplify notation.
@@
Assumption 2 in either form is based on the sequential exchangeability of Robins (1986) in the case where the covariate history consists only of lagged outcomes.222 Other names for Assumption 2b in the literature include “conditional exchangeability”, “sequential ignorability”, and “conditional parallel trends”.
Assumption 2 only restricts a stream of potential outcomes where the first $s-1$ treatment indices conform to the realized history of treatments $D^{t-1}=d^{t-1}$; it imposes conditional independence between present treatment for the entire stream of counterfactual potential outcomes in the future. Mean sequential exchangeability (Assumption 2b) relaxes full sequential exchangeability (Assumption 2a) in two ways. First, the joint independence of treatment with the vector of present and future potential outcomes is replaced with marginal independence of treatment with potential outcomes in each future period. Second, such marginal independence is further relaxed to marginal mean independence. Nonetheless, each version is consistent with the overarching motivation of random assignment of treatment conditional on the observed history.
It is well-known that sequential exchangeability is sufficient for identifying a variety of treatment effects (e.g, Robins and Hernán (2008)). Instead, our focus will be to show how this condition also leads to causal interpretations of alternative estimands in dynamic panel data models for observed outcomes.

### v2

# Heterogeneous Treatment Effects via Linear Dynamic

@@ Abstract
We study the identification of heterogeneous, intertemporal treatment effects (TE) when potential outcomes depend on past treatments. First, applying a dynamic panel data model to observed outcomes, we show that an instrumental variable (IV) version of the estimand in Arellano and Bond (1991) recovers a non-convex (negatively weighted) aggregate of TE plus non-vanishing trends. We then provide conditions on sequential exchangeability (SE) of treatment and on TE heterogeneity that reduce such an IV estimand to a convex (positively weighted) aggregate of TE. Second, even when SE is generically violated, such estimands identify causal parameters when potential outcomes are generated by dynamic panel data models with some homogeneity or mild selection assumptions. Finally, we motivate SE and compare it with parallel trends (PT) in various settings with experimental data (when treatments are sequentially randomized) and observational data (when treatments are dynamic, rational choices under learning).
@@ 1 Introduction
Consider a panel data setting in which a researcher observes individual outcomes $Y_{it}$ and treatments $D_{it}$ over multiple periods t=1,,Tt=1,...,T. At any given time $t$, potential outcomes depend on a full sequence of potential treatments in the past, i.e., $Y_{it}(d^{t})$ where $d^{t}\equiv(d_{s})_{s=1}^{t}$ with $d_{s}\in\{0,1\}$. The objective of causal inference in this context is to learn about certain moments of $Y_{it}(d^{t})$ using the joint distribution of $\{D_{it},Y_{it}\}_{t\leq T}$, where the observed outcomes are determined as $Y_{it}=Y_{it}(D_{i}^{t})$.
We consider the question of causal inference through the lens of linear dynamic panel data models that also include a lagged outcome as a regressor, e.g.,
@@
There are at least two reasons one might argue for the necessity of such a dynamic panel data model (DPDM). First, treatments may only be randomly assigned conditional on the past observed history. For example, when estimating the effect of a training program on earnings or employment outcomes, enrollment in the program may depend on previous earnings or employment (Ashenfelter, 1978). In such cases, the lagged outcome is a necessary control for the exogeneity of the treatment $D_{it}$. Even though (1) appears to deviate from a static panel data model only by an additional control, the fact that this additional control is a lagged dependent variable raises concerns about the strict exogeneity condition on $(D_{it},Y_{it-1})$ that is necessary for two-way fixed-effect (TWFE) regressions. Second, outcomes may exhibit state-dependence. For example, past employment may have a causal effect on present employment (Heckman, 1981). In such cases, the lagged outcome also has a causal effect and may furthermore mediate the effects of past treatment. A DPDM such as (1) allows researchers to disentangle the partial effects of $D_{it}$ from the state dependence in $Y_{it-1}$.
Section 2 introduces the potential outcomes framework underlying our analysis. We then formally state the sequential exchangeability assumption, an identifying condition extensively developed in the biostatistics literature (Robins, 1986), and we relate it to the parallel trends assumption that has been used in the econometrics literature. The sequential exchageability assumption requires that present potential outcomes are (mean) independent from contemporary treatment, once conditioned on a full history of observed outcomes and treatments.
Section 3 investigates whether commonly used dynamic panel data estimators (DPDEs) for observed outcomes — most notably the first-differenced IV (and more generally, GMM) estimators of Anderson and Hsiao (1982) and Arellano and Bond (1991) — can be mapped to causal estimands of interest. These estimators use lagged regressors as instruments to consistently estimate $\beta,\gamma$ in the first-differenced transformation of (1):
@@
where $\Delta$ denotes the first-difference operator. To answer this question, we begin by deriving a general causal decomposition of the estimands of these DPDEs (Lemma 1). Implicit in this decomposition, and more generally in our goal of causal inference, is the possibility that (1) and (2) are not structural models of the outcome-generating process, for example, because of possible heterogeneity in treatment effects (TEs). Our analysis reveals a link between the instrumented first-difference estimators in a reduced-form model of observed outcomes (1) and the heterogeneous TEs on potential outcomes.
@@
This exercise bears resemblance to, but differs qualitatively from, the work of De Chaisemartin and d’Haultfoeuille (2020) and Goodman-Bacon (2021) in decomposing and interpreting TWFE regressions. In a non-dynamic setting where potential outcomes are indexed only by contemporary treatment $Y_{it}(d_{t})$, they showed the TWFE estimand in static panel data models, which is a special case of (1) with $\gamma=0$, is a (possibly negatively) weighted average of heterogeneous TE, $Y_{it}(1)-Y_{it}(0)$.222 The use of TWFE in static panel data models requires that $D_{it}$ be strict exogenous in order to identify $\beta$ and $\gamma$. However, the difference-in-differences approach in De Chaisemartin and d’Haultfoeuille (2020) does not impose that the true data-generating process satisfies the parametric, reduced-form $Y_{it}=\beta D_{it}+\theta_{t}+\alpha_{i}+\varepsilon_{it}$. Their focus is instead on relating the TWFE estimand to aggregated heterogeneous treatment effects under parallel trends; they also consider first-difference regressions under the same assumptions. This literature has also considered the role and the complications of covariates (Heckman, Ichimura, and Todd, 1997; Abadie, 2005; Caetano and Callaway, 2023), as well as proposing alternative estimators under parallel trends (PT) and other assumptions on the outcome or treatment processes, like staggered designs (e.g.,De Chaisemartin and d’Haultfoeuille (2020); Callaway and Sant’Anna (2021); Sun and Abraham (2021)).
In comparison with these earlier works, we face different challenges, because the treatment effects in our setting vary with the full sequence of past potential treatments, and the estimator in question uses instruments designed to address the failure of strict exogeneity of observed $D_{it}$ and $Y_{it-1}$ in the DPDM (1). Among existing works that build on variants of parallel trends, the setting of de Chaisemartin and D’Haultfoeuille (2024) is most similar in that it also allows for intertemporal effects on potential outcomes in non-staggered designs; in contrast, however, our approach is based on sets of new assumptions different than parallel trends.
Building on our decomposition in Lemma 1, we derive sufficient conditions under which an estimand using lagged variables as instruments in the first-differenced equation, as in Anderson and Hsiao (1982) and Arellano and Bond (1991), is guaranteed to be a positively weighted average of heterogeneous TE (Proposition 2). These conditions include the sequential exchangeability assumption of Robins (1986) as well as restrictions on the heterogeneity of TE.
@@
By invoking a condition of sequential exchangeability, our approach is related to, and complements, a broad suite of treatment effect estimators collectively referred to as $g$-methods that have been developed for causal inference in this setting [e.g., see Robins (1986, 1997) and Robins and Hernán (2008) for a survey]. Most closely related, Robins (1997) proposes an IPW estimator that identifies $E[Y_{it}(d^{t})]$ using the observed outcomes $Y_{it}$ in stratum $\{D_{i}^{t}=d^{t}\}$; these potential outcomes can in turn be used to identify history-dependent treatment effects such as $E[Y_{it}(d^{t-1},1)-Y_{it}(d^{t-1},0)]$. We connect our approach back to this existing estimator through an outcome-modifying first-differenced IV estimator (Proposition 3), which recovers the average treatment effect $E[Y_{it}(D^{t-1},1)-Y_{it}(D^{t-1},0)]$ while dispensing with substantive restrictions beyond sequential exchangeability in g-methods.
@@
The sequential exchangeability condition has played a central role in our results above (i.e., causal interpretation using DPDEs, and the outcome-modifying IV estimator). This naturally raises two follow-up questions: (a) Can causal inference be conducted using DPDEs without sequential exchangeability? (b) Can sequential exchangeability be justified in observational data where treatments are endogenously determined by forward-looking individuals? We provide conditions under which the answers to both questions are indeed positive.
@@
To answer (a), Propositions 4, 5, and 6 delineate cases where sequential exchangeability is generically violated, and yet existing or modified DPDEs identify well-defined causal estimands. Specifically, we first focus on a parametric case where the data-generating process of the latent potential outcomes $Y_{it}(d^{t})$ follows a linear DPDM:
$Y_{it}(d^{t})=\beta^{\ast}d_{t}+\gamma^{\ast}Y_{it-1}(d^{t-1})+\theta_{t}^{\ast}+\alpha_{i}^{\ast}+\varepsilon_{it}^{\ast}(d_{t})\text{.}$
@@
We show that in this case, the observed outcomes $Y_{it}=Y_{it}(D_{i}^{t})$ conform with a reduced-form DPDM in (1) in the following sense: (i) the coefficients in the DPDM coincide with constant causal effects, i.e., $\beta=\beta^{\ast},\gamma=\gamma^{\ast}$, and (ii) an assumption of conditional sequential exogeneity with respect to the structural errors $\varepsilon_{it}^{\ast}(\cdot)$ ensures the treatments $D_{it}$ are sequentially orthogonal to the implied reduced-form errors $\varepsilon_{it}\equiv\varepsilon^{*}_{it}(D_{i}^{t})$. Together, (i) and (ii) guarantee the existing DPDEs consistently estimate the constant causal effects, even when sequential exchangeability in Robins (1986) does not hold.333 The conditional sequential exogeneity of contemporaneous treatments and past outcomes in (ii) only holds after controlling for unobserved fixed effects. Hence, it allows for endogenous selection on such unobservables and violates the sequential exchangeability in Robins (1986).
We extend the results for a more general model where only the untreated potential outcomes $Y_{it}(d^{t-1},0)$ take an auto-regressive form above while the intertemporal TEs $\tau_{it}(d^{t-1})\equiv Y_{it}(d^{t-1},1)-Y_{it}(d^{t-1},0)$ are time-varying and heterogeneous. We show that in this case, an IV regression conditional on observed history identifies a conditional average treatment effect on the treated, as well as time trends and the state dependence parameter $\gamma^{*}$. This result requires two mild assumptions: conditional serial uncorrelation of the structural errors of potential outcomes $\varepsilon^{*}_{t}(d^{t-1})$ (Assumption 6), and conditional mean independence of heterogeneous TE from past errors $\varepsilon_{is}^{*}(\cdot)\text{ for }s\leq t-1$ (Assumption 7). The latter is a weak restriction on the heterogeneous TEs in that it does not impose any “conditional unconfoundedness” condition. That is, it allows for conditional correlation between TEs and contemporary treatment choices.
To investigate question (b), Section 4 provides, in the context of both experimental and observational data, conditions under which sequential exchangeability holds whereas a common alternative of parallel trends does not. For an experimental setting, we show sequentially randomized treatments are compatible with sequential exchangeability, but not compatible with parallel trends in general.
@@
We also provide critical guidance for empiricists to assess sequential exchangeability in observational data, especially when treatment choices are most likely dependent upon past outcomes. For example, when estimating the effect of a training program on earnings or employment outcomes, enrollment into the training program may depend on previous earnings and employment, as per Ashenfelter (1978).
@@
To this end, we consider sequential exchangeability in a model of dynamic choices that serves as a structural motivation. A particular focus of the model is on the possibility of learning, which is compatible with the nature of sequential exchangeability assumptions that condition (like a learning decision-maker) on the history of observed outcomes. The structural model provides a coherent dynamic decision problem of an economic agent where preferences, beliefs and decision rules are specified and shows the kinds of restrictions that are required, especially on the distribution of prior beliefs, to yield the sequential exchangeability conditions we use. Thus, our work connects to recent work on identification of learning models in nonparametric settings (e.g., (Bunting, Diegert, and Maurel, 2022)). This exercise is similar to the one done in Marx, Tamer, and Tang (2024) for a model under parallel trends.
@@
Our work also relates to a broader agenda relating sequential exchangeability and parallel trends (e.g., Renson, Hudgens, Keil, Zivich, and Aiello (2023)) and the corresponding estimators, for example through “bracketing” relationships on the target parameter (Angrist and Pischke, 2008; Ding and Li, 2019). Additionally, Han (2021) studies the identification of dynamic treatment effects beyond sequential exchangeability, namely under a dynamic version of rank similarity. In complementary work, Kim (2023) studies identification of a dynamic panel data model when treatment is staggered and potential outcomes are non-negative, statically indexed, and take a multiplicative form; in this case, he obtains identification results for ratios of treatment effects under sequential exchangeability conditional on persistent heterogeneity. Finally, Klosin (2024) provides an interesting approach that corrects for the bias caused by ignoring dynamic feedback in static panel data regressions.
The rest of the paper is organized as follows. Section 2 introduces the dynamic potential outcome framework. Section 3 studies how to use DPDEs to make causal inference about intertemporal heterogeneous TEs in this framework, with and without the sequential exchangeability condition. Section 4 investigates when sequential exchangeability can be justified in an (observational) context of dynamic treatment choices with learning, and an (experimental) setting of sequential randomization of treatments. Section 5 concludes. Proofs are collected in Appendix A.
@@ 2 A Dynamic Framework for Potential Outcomes
The researcher observes treatment $D_{it}$ with realized outcomes $Y_{it}$ for individual units $i$ in time periods $t=0,1,\dots,T$. There is a pre-treatment period $t=0$ with an initial realized outcome $Y_{i0}$ and $\Pr\{D_{i0}=0\}=1$. We do not model how $Y_{i0}$ is determined prior to the sampling period, but allow it to be correlated with potential outcomes and treatment choices in our analysis. Potential outcomes in period $t\geq 1$ are functions only of an individual’s own treatment history through period $t$.
@@
Potential outcomes for individual unit $i$ in period $t$ are indexed only by $i$’s own treatment through that time period, i.e., $Y_{it}(d_{i}^{t})$, with $d^{t}_{i}\equiv(d_{is})_{s=1}^{t}$.
Assumption 1 rules out spillovers across individual units as well as anticipation of future treatments affecting present potential outcomes. Observed outcomes equal potential outcomes evaluated at observed treatments, $Y_{it}=Y_{it}(D_{i}^{t})$. In this framework, the treatment effects $Y_{it}(d^{t})-Y_{it}(\widetilde{d}^{t})$ are intertemporal in that they are defined by the full history of past potential treatments.
We adopt notation $Y_{i}^{t}\equiv(Y_{i0},Y_{i1},\dots,Y_{it})$ for the history of realized outcomes, $Y_{i}^{t}(d^{t})\equiv(Y_{i0},Y_{i1}(d^{1}),\dots,Y_{it}(d^{t}))$ for that of potential random outcomes, and $Y_{i}^{t}(\cdot)\equiv\{Y_{i}^{t}(d^{t}):d^{t}\in\{0,1\}^{t}\}$, i.e., the collection of potential random vectors through period $t$. Throughout, we restrict to realized and potential treatment vectors $D_{i}^{t}$ and $d_{i}^{t}$ with degenerate initial realization $D_{i0}=d_{i0}=0$, which we suppress where convenient.
The main concern that motivates the inclusion of lagged outcomes as controls in a dynamic panel data model (DPDM) for observed outcomes in (1) is that the potential outcomes are (mean) independent from the treatments only after conditioning on the observed history of past treatments and outcomes. The assumption below formalizes such conditional independence. For subsequent references, we include both a full and mean version of the assumption. We drop individual subscripts $i$ to simplify notation.
@@
Assumption 2 in either form is based on the sequential exchangeability condition in Robins (1986) when the covariate history consists only of lagged outcomes.444 Other names for Assumption 2b in the literature include “conditional exchangeability”, “sequential ignorability”, and “conditional parallel trends”.
Assumption 2 only restricts a stream of potential outcomes where the first s1s-1 treatment indices conform to the realized history of treatments $D^{t-1}=d^{t-1}$; it imposes conditional independence between present treatment and the entire stream of counterfactual potential outcomes in the future. Mean sequential exchangeability (Assumption 2b) relaxes full sequential exchangeability (Assumption 2a) in two ways. First, the joint independence of treatment with the vector of present and future potential outcomes is replaced with marginal independence of treatment with potential outcomes in each future period. Second, such marginal independence is further relaxed to marginal mean independence. Nonetheless, each version is consistent with the overarching motivation of random assignment of treatment conditional on the observed history.
It is well-known that sequential exchangeability is sufficient for identifying a variety of treatment effects (e.g, Robins and Hernán (2008)). Instead, our focus will be to show how this condition also leads to causal interpretations of alternative estimands in the dynamic panel data models for observed outcomes.